RC Update and Run plan for Sep 15,2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are M-F at 4:15 PM

To join the Meeting:  https://bluejeans.com/861439920
To join via phone :  +1.408.740.7256 (US (San Jose)), +1.888.240.2560 (US Toll Free))

Tuesday 15 Sept 2020.

General updates

(See d2n Run Plan 2020 for broader status/overview.)

  • Target work during beam studies yesterday was successful
    • Replaced spectrum analyzer mounting plate. [1], [2], [3]
    • Measured target field direction. [4] These measurements are consistent with Bill/Murchhana's earlier measurements. So target field angle is not the driver for lower than expected polarization.
    • Rotated target to Transverse/270 -- Done [5], [6]
    • Took some parasitic Carbon-Optics runs during CW portion of beam studies in Kin-A and Kin-X
  • Target Oven temperature increased from 205 to 208 [7]
  • Beam handed back to us in early OWL, and continued production running in Kin-A, Kin-X at transverse/270
    • Alternating between triggers T1 (1hr), T2 (3hrs); then repeating
  • Lost SL 2K coldbox at 8am Tuesday morning [8]; minimum 4 hour down, cryo identified the problem but still working on repair (as of 9:15am)
    • Doing QWP target optimization during this down [9]
    • NMR @ ~42% with calibration constant 5.04 (1:13pm) [10] (Note: calibration constant is estimated).
    • Still a few more hours (as of 1:30pm) [11]
  • Will continue production running in Kin-A, Kin-X at transverse/270 (w/ T1 and T2 triggers) after standard beam profile checks.

Notes for the Future

  • Moller measurement planned for Wednesday DAY starting at 8:30am
    • Hall A's Moller will run Wednesday DAY as well. Will need to coordinate on IHWP switches.
  • Melanie, Murchhana, Chao, Sylvester: please check good electron yields per Coulomb for Kin-A/X runs with T1 vs. T3 (T2)
    • Mostly complete. Early results don't show any surprises..
  • EPR calibrations cross checks are ongoing [Melanie]
    • So far it all looks good.
    • Summary log entries: EPR(270, Kin-Z) [12], EPR(90, Kin-X) [13]
  • Bill, Murchhana check compass/field measurements
    • Also checking with Jixie re: modeled fields
    • Need to make a write-up that clearly defines all the different coordinate systems used, swapped power supply/polarity issues, etc.
  • 1-pass running transition scheduled for 6am Friday, Sept 18
    • RadCon is aware of the plans
    • Hall A will setup the charge-asym feedback system for us (P. King, et al)

Pending Issues

  • Review and do any final prep/planning for 1-pass running
    • Paddle optimization for 3He Elastic xsec running [Chao?]
    • EPR calibration for 1-pass 8.5deg setting (estimate from prior calibration?)
  • Beam energy drifts
    • Should confirm that MCC remembers to shift the energy lock to Hall C when Hall A turns off
    • Just something to keep an eye on. Has not been a problem lately.
    • Hall C 5-pass polarization goes down as energy (HALLC:p) goes up [14]

Opportunistic Jobs/Access

Current Run Plan

Tuesday DAY → Tuesday SWING

  • Wait for 2K coldbox repair, and south linac to be restored
  • Continue production data in Kin-A, Kin-X with pol 3He target, Transverse/270 polarization with pNMR/NMR every 5 hrs.
    • NOTE: Take 1 (1hr) run with T1(3/4) trigger; then 3 runs (1hr each) using T2(EL_real) trigger; Then repeat that cycle.
      • Be sure to monitor and set the prescales correctly.
      • Put T1 or T2 in the CODA comments field for the runs.
      • Note the T1 or T2 trigger type for each run in the Shift Summary and Run database using the add_run_comment command.

Wednesday OWL → Wednesday SWING

  • Moller measurement on Wednesday DAY at 8:30am (Murchanna, Bill)
    • This will also be used to infer polarization at 1-pass (if it fails, we'll try again on Thursday DAY)
  • Wednesday DAY during Moller Move spectrometers to Kin-C, Kin-Z and continue Transverse/270 running there
  • Continue production data in Kin-C, Kin-Z with pol 3He target, Transverse/270 polarization with pNMR/NMR every 5 hrs when Moller complete.
    • All runs for Kin-C and Kin-Z settings use the T1 trigger only

Thursday OWL → Thursday SWING

  • Continue production data in Kin-C, Kin-Z with pol 3He target, Transverse/270 polarization with pNMR/NMR every 5 hrs.
    • All runs for Kin-C and Kin-Z settings use the T1 trigger only

Friday OWL

  • Continue production data in Kin-C, Kin-Z with pol 3He target, Transverse/270 polarization with pNMR/NMR every 5 hrs.
    • All runs for Kin-C and Kin-Z settings use the T1 trigger only
  • 6 am Move to Controlled Access (w/ RadCon support) to begin transition to 1-pass

Friday DAY

  • (Started at 6am) Open target enclosure
    • Install collimators (JP, RadCon)
    • Check on EPR photo-diode status, likely replace it? (JP, RadCon)
  • Ensure fast raster is reconfigured correctly (Expert: Jones, Gunning)
  • Rotate to minimum angles for HMS (11.7) and SHMS (8.5) (S. Lassiter, A. Comer)
  • Configure target for Longitudinal/0 or Longitudinal/180 (TBD)
  • Close up and wait for MCC to establish (1-pass) beam to the Hall

Friday SWING → Monday

  • See task list at Run Plan for 1-pass for more details
    • Initial priority will be PbPt polarization cross check with longitudinal pol. elastics.
    • These have lower priority but will likely be pursued while target spins up
      • 3He, N2 density measurements
      • H2 xsec measurement
      • Pressure curves


Standing Orders

  • Change the IHWP at least every 8 hours when running on glass cells. Call Halls A and B in advance. If there are any issues, call the RC
  • Ensure that all information in the CODA run-start GUIs is correct for every run!
  • If beam goes away for more than 3h, or MCC does operations on the beam that may change the beam tune, then move target to No Target position,
    • verify Harps scans at 3H07A and 3H07B are within spec, and
    • ensure beam is stable (with all locks ON) at 0, 0mm before continuing.
    • Beam recovery procedure has more detailed procedure if needed.
  • New beam positions have been verified and established to be (0,0) mm for all targets.
    • Monitor that beam is at (0,0) using the stripchart on the Wall screen and verify no scraping on the glass cells using the Analysis monitor.
  • Check the position locks are on.
  • Check that CODA is running normally
  • Check online replays shortly after starting runs, & full replay results too.
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Follow the updated guidelines in the How-to#Spectrometer_Magnets Spectrometer Magnets How to
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • If any magnet trips, snapshot the associated power supply window and post it to hclog before resetting the interlock.
  • Monitor the beam position on target and call MCC immediately if you see large excursions (1 mm level).
  • When rotating the spectrometer, use HCCAM01 to keep an eye on the SOLID Cherenkov test stand when rotating SHMS (in addition to watching the live angle-cam image, etc). The updated rotation instructions are here: https://hallcweb.jlab.org/wiki/index.php/Spectrometer_Rotation#NOTICES_IN_EFFECT

Useful Links

PolHe3 Polarization Measurements

Misc Links